# Ensemble-Based Event Camera Place Recognition Under Varying Illumination

### v1

@@ Abstract
Compared to conventional cameras, event cameras provide a high dynamic range and low latency, offering greater robustness to rapid motion and challenging lighting conditions. Although the potential of event cameras for visual place recognition (VPR) has been established, the development of robust VPR frameworks under severe illumination changes remains an open research problem. In this paper, we introduce an ensemble-based approach to event camera place recognition that combines sequence-matched results from multiple event-to-frame reconstructions, VPR feature extractors, and temporal resolutions. Unlike previous event-based ensemble methods limited to temporal resolution, our broader fusion strategy delivers significantly improved robustness under varied lighting conditions (afternoon, sunset, night, etc.), achieving 57% relative improvement in Recall@1 across day-night transitions. We evaluate our approach on two long-term driving datasets (with 8 km per traverse) without metric subsampling, thereby preserving natural variations in speed and stop duration that influence event density. We also conduct a comprehensive analysis of key design choices, including binning strategies, polarity handling, reconstruction methods, and feature extractors, to identify the most critical components for robust performance. Additionally, we propose a modification to the standard sequence matching framework that enhances performance at longer sequence lengths. To facilitate future research, we will release our codebase and benchmarking framework.
Figure 1: Overview of our ensemble-based event camera place recognition pipeline. (A) Raw event data $(x,y,t,p)$ is first segmented into discrete bins using a fixed time duration. (B) Each bin is reconstructed into a 2D frame using one of several methods: event count (with or without polarity), time surface [1], or the learned E2VID model [2]. (C) The reconstructed frames are passed through visual place recognition (VPR) feature extractors, including NetVLAD [3], CosPlace [4], MixVPR [5], and MegaLoc [6], to generate global descriptors. (D) Pairwise descriptor similarities are computed, and sequence matching is applied to each resulting similarity matrix. Finally, sequence scores are aggregated across multiple reconstruction methods, feature extractors and temporal resolutions to enhance recognition performance.
@@ I Introduction
Event cameras are a novel class of vision sensors that record changes in brightness asynchronously and independently for each pixel, producing a stream of events that encode the time, pixel location, and polarity of intensity changes. Unlike conventional frame-based cameras, event cameras offer microsecond temporal resolution, high dynamic range, and low latency, making them particularly suitable for high-speed, low-power applications and operation in challenging lighting conditions [7, 8, 9].
These characteristics have generated growing interest in applying event cameras to autonomous navigation tasks [10, 11, 12, 13, 14, 8], including visual place recognition (VPR), a core component of long-term localisation and map-based re-localisation. Traditional VPR systems rely on frame-based imagery and aim to extract features invariant to appearance changes, dynamic elements, viewpoint shifts, and lighting variation [15, 16, 17, 18, 19]. However, they often falter under extreme illumination changes (e.g., day-night)[20], where event cameras can still provide a useful signal. Despite their promise, event-based VPR methods remain underexplored, especially in conditions involving severe illumination variation. Key open research questions include: 1) how to best represent and reconstruct event data for downstream place recognition, and 2) whether combining complementary representations or VPR models can improve robustness under challenging conditions.
We present an ensemble-based approach to event camera place recognition that enhances robustness through late score fusion. Specifically, we propose three ensemble strategies: (i) aggregating across different event-to-frame reconstruction methods, both classical (e.g., event count, time surface [1]) and learned (E2VID [2]); (ii) aggregating across multiple state-of-the-art (SoTA) VPR feature extractors, including NetVLAD [3], CosPlace [4], MixVPR [5], and MegaLoc [6], and (iii) aggregating all configurations with temporal resolutions, event-to-frame reconstruction methods and VPR feature extractors. Unlike prior event-based VPR ensembling that fuses predictions only across temporal resolutions [11], our strategy evaluates ensembles of reconstruction methods, feature extractors and temporal resolutions, yielding improved robustness under challenging appearance changes, including illumination shifts across the day. We also use complete traverses without metric subsampling in our evaluations, which preserves natural variations in speed and stop duration that are challenging in event-based settings where event density is motion dependent. Finally, we apply a modified sequence matching strategy to each ensemble member, leveraging temporal consistency for improved performance.
@@
    An ensemble-based event VPR framework that aggregates predictions across reconstruction methods, feature extractors and temporal resolutions, achieving an average 17% absolute improvement and a 57% relative gain in Recall@1 under day–night transitions.
@@ II Related Works
II-A Visual Place Recognition
Visual place recognition (VPR) identifies the best match for a query image within a reference database and serves as a vision-based global re-localisation method for navigation systems [15, 16, 17, 18, 19]. It typically involves extracting feature vectors from images and computing pairwise similarity to find the closest match.
Modern VPR methods using conventional frame-based cameras rely heavily on deep learning. NetVLAD [3] was an early breakthrough, combining a VGG-16 backbone with a VLAD (Vector of Locally Aggregated Descriptors) layer for end-to-end place representation. Subsequent methods have aimed to improve scalability and performance. CosPlace [4] improves scalability by using a ResNet-50 backbone and training on spatially and directionally disjoint classes to extract distinctive global descriptors. MixVPR [5] advances descriptor learning by applying MLP-based mixing to flattened ResNet-50 feature maps, enabling holistic aggregation with reduced compute. The current state-of-the-art, MegaLoc [6], employs a DINOv2-Base ViT backbone with a SALAD aggregation layer, achieving robust performance across VPR, landmark retrieval, and visual localisation tasks.
In our work, we adopt all four of these methods, NetVLAD, CosPlace, MixVPR and MegaLoc, for event-based place recognition, evaluating both their performance and the benefits of ensembling these feature descriptors.
@@ II-B Event-Based VPR
Event cameras output a continuous stream of sparse, asynchronous events rather than conventional image frames. To leverage standard vision pipelines, particularly those based on deep learning, these events are typically aggregated into frame-like representations or voxel tensors.
An early study on event-based visual SLAM [24] demonstrated the feasibility of place recognition over a 2.7km route using SeqSLAM [21] applied to polarity-removed event count reconstructions. Fischer et al.[11] later introduced an ensemble-based approach that fused multiple temporal windows, reconstructing event frames with E2VID[2] and extracting features using NetVLAD. In follow-up work [13], they proposed a lightweight method using a sparse subset of varying pixels as descriptors and sum of absolute differences (SAD) for matching, achieving competitive performance with minimal compute cost.
@@
Although Ensemble-Event-VPR [11] and Event-VPR [14] conducted ablations on windowing strategies, reconstruction methods, network backbones and loss functions, several aspects remain underexplored. In particular, prior work offers limited comparison of time surface reconstructions, polarity handling strategies, event binning methods, and the performance of various SoTA VPR feature extractors when applied to event data, especially under day–night transitions. Our work addresses these gaps through an evaluation of these design decisions and demonstrates their significant impact on VPR performance under varying illumination conditions.
Other efforts explored hardware optimisations and deployment. One study applied a closed-loop bias controller for brightness-adaptive VPR, using SAD matching on event count reconstructions[26]. Most recently, Hines et al. [8] demonstrated event-based place recognition using a spike-based encoder with event count frames accumulated over one second, deployed on ultra-low-power neuromorphic hardware (SynSense Speck).
@@ II-C Ensemble Methods
Ensembles improve prediction accuracy when individual models are both accurate and diverse [27]. If individual models have uncorrelated errors and perform better than random guessing, the probability that a majority vote is incorrect becomes lower than the individual error rate [28]. As a result, ensembling is a core strategy in many machine learning approaches, including methods such as bagging [29], boosting [30], and mixture of experts [31]. These ensemble techniques have been successfully applied across a wide range of domains, from classification tasks to large-scale language and vision models.
Ensemble methods have also been applied across various stages of the VPR pipeline to improve robustness and generalisation. Hierarchical Multi-Process Fusion [32] performs late-stage ensemble re-ranking by filtering candidate matches across multiple pipelines. Patch-NetVLAD [33] aggregates local patch features at multiple scales to improve viewpoint robustness. In neuromorphic settings, an ensemble of spiking neural networks [34] independently learns non-overlapping regions and combines spike responses from all experts at inference, suppressing hyperactive neurons to improve global matching. For event-based VPR, Fischer et al. [11] ensemble predictions across multiple time resolutions. Our approach builds on these ideas by ensembling across reconstruction methods, feature extractors and temporal resolutions, incorporating late-score fusion in a unified benchmarking framework.
@@ III-A Event Slicing
The event camera records a continuous stream of events where each event $e_{j}=(x_{j},y_{j},t_{j},p_{j})$ contains the location of the pixel $(x_{j},y_{j})$, the timestamp $t_{j}$, and the polarity $p_{j}\in\{-1,+1\}$, representing a positive or negative brightness change.
@@ III-B Event to Frame Reconstruction
The first method is a learned reconstruction model, E2VID [2], which produces an image estimate from a sequence of events using an encoder-decoder architecture with spatiotemporal fusion. We denote the reconstructed frame from bin $\mathbf{B}_{i}$ as
$\mathbf{I}^{\text{E2VID}}_{i}=\mathcal{R}_{\text{E2VID}}(\mathbf{B}_{i}),$ (3)
where $\mathcal{R}_{\text{E2VID}}$ is the pretrained E2VID model applied to the events in bin $\mathbf{B}_{i}$.
@@ III-C Feature Extraction and Descriptor Matching
Let $\mathbf{I}^{Q}_{q}$ and $\mathbf{I}^{R}_{j}$ denote the $q$-th query frame and the $j$-th reference frame, respectively. A descriptor extractor $f(\cdot)$ maps each frame to a descriptor vector in a common embedding space:
$\mathbf{d}^{Q}_{q}=f(\mathbf{I}^{Q}_{q}),\quad\mathbf{d}^{R}_{j}=f(\mathbf{I}^{R}_{j}),\quad\mathbf{d}\in\mathbb{R}^{D},$ (6)
where $D$ is the descriptor dimension, $Q$ is the size of the query set and $R$ is the size of the reference set. We compute the similarity between each query and reference descriptor using the negative L2 distance, and define the similarity matrix $\mathbf{S}\in\mathbb{R}^{Q\times R}$ as:
@@ III-D Sequence Matching
To improve robustness near sequence boundaries and mitigate local biases caused by visual aliasing, we extend SeqSLAM by introducing a variable-length matching kernel and dual-axis normalisation. While SeqSLAM computes per-row diagonal scores using a fixed-size convolutional kernel, our method adapts the effective kernel size near sequence boundaries. Let $R$ denote the number of reference frames and $L$ the desired sequence length. For each query index $q$, we construct a submatrix $\mathbf{C}$ of the similarity matrix $\mathbf{S}\in\mathbb{R}^{Q\times R}$ by extracting up to $L$ preceding query frames (the available history in a real time deployment setting), without exceeding the matrix boundary:
$\mathbf{C}_{q}\in\mathbb{R}^{N\times R},\quad\text{where}\quad N=\min(L,q+1).$
We apply z-score normalisation to $\mathbf{C}$ across columns and then across rows, to reduce local bias in the similarity submatrix. To compute the sequence-matched score at each reference index $j$, we calculate the trace of the most recent $k\times k$ diagonal block:
$\tilde{\mathbf{S}}[q,j]=\operatorname{trace}\left(\mathbf{C}_{[N-k:N,\;j-k+1:j+1]}\right),\quad k=\min(N,j+1).$
@@ III-E Ensembling Strategies
Each $\tilde{\mathbf{S}}^{(m)}\in\mathbb{R}^{Q\times R}$ is the sequence-matched similarity matrix produced by the $m$-th method configuration. We treat these as individual ensemble members:
$\mathcal{S}=\left\{\tilde{\mathbf{S}}^{(1)},\ \dots,\ \tilde{\mathbf{S}}^{(M)}\right\}.$ (8)
To aggregate predictions across configurations, we perform element-wise summation of all aligned similarity matrices, followed by maximum similarity selection:
@@ IV-A Datasets
In this paper, we evaluate our method on two outdoor datasets: BrisbaneEventVPR[11] and NSAVP[23].
Brisbane Event VPR is a widely used benchmark dataset for event-based visual place recognition, introduced in 2020 and evaluated in several works [13, 10, 12, 8]. It consists of six traverses of an 8km route in Brisbane, Australia, recorded under various weather and illumination conditions, including daytime, night, sunrise, sunset, and morning on single and multilane roads. Data was captured using a DAVIS 346 event camera mounted on the vehicle’s windshield, along with GPS and 1080p RGB video from a consumer-grade camera.
Novel Sensors for Autonomous Vehicle Perception (NSAVP) is a more recent dataset collected in Michigan, USA, with a focus on evaluating modern sensor technologies in diverse and realistic environments. It comprises two distinct routes (8.3 km and 8.6 km) traversed in both directions under varying lighting conditions. Data was collected using a high-resolution (0.31 MP) Prophesee Gen4 event sensor and high-precision 200 Hz ground-truth from an Applanix POS-LV 420 system, mounted as part of a roof-mounted stereo sensor suite. Compared to BrisbaneEventVPR, NSAVP introduces greater environmental diversity, including suburban neighbourhoods and dense urban scenes with multi-story buildings, and offers higher fidelity sensing for large-scale benchmarking. Unlike the benchmark protocol from this dataset [23], which evaluates only 6.7 km of the 8.3 km route with frames sampled every 2 meters, we use the full traverses without metric subsampling. This preserves natural variations in speed and stop duration, which are critical in event-based settings where event density is motion-dependent.
We use five day-day and five day-night conditions to enable fair comparison across illumination changes. For BrisbaneEventVPR [11], we use sunset1 as the reference and query against sunrise, morning, daytime, sunset2 and night. For NSAVP [23], we evaluate five pairs from Route 0: forward-afternoon vs. forward-sunset, forward-afternoon vs. forward-night, forward-night vs. forward-sunset, reverse-afternoon vs. reverse-night, and reverse-sunset vs. reverse-night. We follow the naming convention from the original paper, where R0 denotes Route 0, F/R indicate forward/reverse traversal, and A, S, N correspond to afternoon, sunset, and night conditions, respectively. Other datasets, such as DDD20 [35] and NYC Events [36], are not considered in this work, as they lack explicit repeat traversals.

### v2

@@ Abstract
Compared to conventional cameras, event cameras provide a high dynamic range and low latency, offering greater robustness to rapid motion and challenging lighting conditions. Although the potential of event cameras for visual place recognition (VPR) has been established, developing robust VPR frameworks under severe illumination changes remains an open research problem. Here, we introduce an ensemble-based approach to event camera place recognition that combines sequence-matched results from multiple event-to-frame reconstructions, VPR feature extractors, and temporal resolutions. Unlike previous event-based ensemble methods, which only utilise temporal resolution, our broader fusion strategy delivers significantly improved robustness under varied lighting conditions (e.g., afternoon, sunset, night), achieving up to 77% relative improvement in Recall@1 across day-night transitions. We evaluate our approach on two long-term driving datasets (with 8 km per traverse) without metric subsampling, thereby preserving natural variations in speed and stop duration that influence event density. We also conduct a comprehensive analysis of key design choices, including binning strategies, reconstruction methods, and feature extractors, to identify the most critical components for robust performance. Additionally, we propose a modification to the standard sequence matching framework that enhances performance at longer sequence lengths. To facilitate future research, we release our codebase and benchmarking framework [^1] https://github.com/theresejoseph/ensemble_event_vpr_bench.
Figure 1: Overview of our ensemble-based event camera place recognition pipeline. A) Each bin is reconstructed into a 2D frame using one of several methods: event count (with/without polarity), time surface [1], or the learned E2VID model [2]. (B) These reconstructed frames are processed by an ensemble of visual place recognition (VPR) feature extractors (NetVLAD [3], CosPlace [4], MixVPR [5], MegaLoc [6]) to generate global descriptors. (C) Pairwise descriptor similarities are computed, followed by sequence matching on each resulting similarity matrix. (D) Finally, sequence scores are aggregated across the multiple reconstruction methods, feature extractors, and temporal resolutions to enhance recognition performance.
@@ I Introduction
Event cameras are a class of vision sensors that record changes in brightness asynchronously and independently for each pixel, producing a stream of events that encode the time, pixel location, and polarity of intensity changes. Unlike conventional cameras, they offer microsecond temporal resolution, high dynamic range, and low latency, making them particularly suitable for high-speed, low-power applications and operation in challenging lighting conditions [7, 8, 9].
These characteristics have generated growing interest in applying event cameras to autonomous navigation tasks [10, 11, 12, 13, 14, 8], including visual place recognition (VPR), a core component of long-term localisation and map-based re-localisation. Traditional VPR systems rely on frame-based imagery and aim to extract features invariant to appearance changes, dynamic elements, viewpoint shifts, and lighting variation [15, 16, 17, 18, 19]. However, they often falter under extreme illumination changes (e.g., day-night)[20], where event cameras can still provide a useful signal. Despite their promise, event-based VPR remains underexplored, especially in conditions involving severe illumination variation. Key open research questions include: 1) how to best represent and reconstruct event data for downstream place recognition, and 2) whether combining complementary representations or VPR models can improve robustness under challenging conditions.
We present an ensemble-based approach to event camera VPR that enhances robustness through late score fusion. Specifically, we propose three ensemble strategies: (i) aggregating across different event-to-frame reconstruction methods, both classical (e.g., event count, time surface [1]) and learned (E2VID [2]); (ii) aggregating across multiple VPR feature extractors, including NetVLAD [3], CosPlace [4], MixVPR [5], and MegaLoc [6], and (iii) aggregating all configurations with temporal resolutions, event-to-frame reconstruction methods and VPR feature extractors.
Unlike prior event-based VPR ensembling that fuses predictions only across temporal resolutions [11], our strategy evaluates ensembles of reconstruction methods, feature extractors and temporal resolutions, yielding improved robustness under challenging appearance changes, including illumination shifts across the day. This broader fusion combines complementary cues from different representations, leading to greater robustness across diverse visual and motion conditions. We also use complete traverses without metric subsampling in our evaluations, which preserves natural variations in speed that are challenging in event-based settings where event density is motion-dependent. Finally, we apply a modified sequence matching strategy to each ensemble member, leveraging temporal consistency for improved performance.
@@
    An ensemble-based event VPR framework that aggregates predictions across reconstruction methods, feature extractors and temporal resolutions, achieving up to 77% relative gain in Recall@1 under day–night transitions.
@@ II Related Works
II-A Place Recognition
Place recognition matches a query to a reference database via feature extraction and similarity search, enabling global re-localisation and loop closure [15, 16, 17, 18, 19, 24]. Visual place recognition (VPR) uses vision-based sensors such as conventional frame or event cameras. Beyond vision, LiDAR-based methods learn or extract geometry-aware point cloud descriptors [25], radar-based methods explore robustness to weather and illumination [26], and multimodal methods combine camera frames with LiDAR or events [27, 28, 29].
Modern VPR relies on deep networks: NetVLAD (VGG-16 + VLAD) [3], CosPlace (ResNet-50 with disjoint classes) [4], MixVPR (MLP mixing on flattened features) [5], and MegaLoc (DINOv2-Base + SALAD) [6]. In our work, we adopt these methods for event-based place recognition, evaluating both their performance and the benefits of ensembling.
@@ II-B Event-Based VPR
Event cameras output a continuous stream of sparse, asynchronous events rather than conventional image frames. Compared to conventional frame-based cameras, they provide microsecond-level temporal resolution, very low latency, high dynamic range, and low power consumption, which makes them well-suited to mobile sensing under fast motion and challenging illumination conditions such as night driving or headlight glare. These sensing properties are directly beneficial for visual place recognition, where robustness to motion blur, abrupt lighting changes, and low-light scenes is essential for reliable matching along long vehicle traverses [30, 31].
To leverage standard vision pipelines, particularly those based on deep learning, these events are typically aggregated into frame-like representations or voxel tensors. An early study on event-based visual SLAM [32] demonstrated the feasibility of place recognition over a 2.7km route using SeqSLAM [21] applied to polarity-removed event count reconstructions. Fischer et al.[11] later introduced an ensemble-based approach that fused multiple temporal windows, reconstructing event frames with E2VID[2] and extracting features using NetVLAD. In follow-up work [13], they proposed a lightweight method using a sparse subset of varying pixels as descriptors and sum of absolute differences (SAD) for matching, achieving competitive performance with minimal compute cost.
@@
Although Ensemble-Event-VPR [11] and Event-VPR [14] conducted ablations on windowing strategies, reconstruction methods, network backbones and loss functions, several aspects remain underexplored. In particular, prior work offers limited comparison of time surface reconstructions, event binning methods, and the performance of various SoTA VPR feature extractors when applied to event data, especially under day–night transitions. Our work addresses these gaps through an evaluation of these design decisions, demonstrating their significant impact on VPR performance under varying illumination conditions. Beyond analysing these factors individually, we show that combining them through score-level fusion further enhances robustness, as complementary representations contribute distinct but reinforcing cues under appearance change.
Other efforts explored hardware optimisations and deployment. One study applied a closed-loop bias controller for brightness-adaptive VPR, using SAD matching on event count reconstructions[34]. Most recently, Hines et al. [8] demonstrated event-based place recognition using a spike-based encoder with event count frames accumulated over a second, deployed on ultra-low-power neuromorphic hardware.
@@ II-C Ensemble Methods
Ensembles improve accuracy when base models are both accurate and diverse. If individual models have uncorrelated errors and perform better than random guessing, the probability that a majority vote is incorrect becomes lower than the individual error rate [35, 36].
In VPR, ensembles enhance robustness at different stages of the pipeline. Hierarchical Multi-Process Fusion applies late re-ranking across pipelines [37], while Patch-NetVLAD aggregates multi-scale patch features for viewpoint changes [38]. In neuromorphic settings, ensembles of spiking networks learn non-overlapping regions and combine spike responses at inference [39]. For event-based VPR, ensembling across time resolutions has shown gains [11]. This method identified that fusion based on the mean of similarity or difference matrices yielded the highest performance across product, median, min, max, trimmed mean and weighted strategies.
We extend these ideas by ensembling across reconstruction methods, feature extractors, and temporal resolutions, by using late score fusion to combine the varied similarity responses of these heterogeneous configurations. This extension enables complementary modalities to reinforce each other, improving matching consistency across changes in lighting, texture, and motion dynamics.
@@ III-A Event Slicing
Event cameras record a continuous stream of events where each event $e_{j}=(x_{j},y_{j},t_{j},p_{j})$ contains the pixel location $(x_{j},y_{j})$, the timestamp $t_{j}$, and the polarity $p_{j}\in\{-1,+1\}$, representing a positive or negative brightness change.
@@ III-B Event to Frame Reconstruction
The first method is a learned reconstruction model, E2VID [2], which produces an image estimate from a sequence of events using an encoder-decoder architecture with spatiotemporal fusion. We denote the reconstructed frame from bin $\mathbf{B}_{i}$ as $\mathbf{I}^{\text{E2VID}}_{i}=\mathcal{R}_{\text{E2VID}}(\mathbf{B}_{i})$, where $\mathcal{R}_{\text{E2VID}}$ is the pretrained E2VID model applied to the events in bin $\mathbf{B}_{i}$.
@@ III-C Feature Extraction and Descriptor Matching
Let $\mathbf{I}^{Q}_{q}$ and $\mathbf{I}^{R}_{j}$ denote the $q$-th query frame and the $j$-th reference frame, respectively. A descriptor extractor $f(\cdot)$ maps each frame to a descriptor vector in a common embedding space: $\mathbf{d}^{Q}_{q}=f(\mathbf{I}^{Q}_{q})$, $\mathbf{d}^{R}_{j}=f(\mathbf{I}^{R}_{j})$, $\mathbf{d}\in\mathbb{R}^{D}$, where $D$ is the descriptor dimension, $Q$ is the size of the query set and $R$ is the size of the reference set. We compute the similarity between each query and reference descriptor using the negative L2 distance, and define the similarity matrix $\mathbf{S}\in\mathbb{R}^{Q\times R}$ as:
@@ III-D Sequence Matching
To improve robustness near sequence boundaries and mitigate local biases caused by visual aliasing, we extend SeqSLAM by introducing a variable-length matching kernel and dual-axis normalisation. While SeqSLAM computes per-row diagonal scores using a fixed-size convolutional kernel, our method adapts the effective kernel size near sequence boundaries. Let $R$ denote the number of reference frames and $L$ the desired sequence length. For each query index $q$, we construct a submatrix $\mathbf{C}$ of the similarity matrix $\mathbf{S}\in\mathbb{R}^{Q\times R}$ by extracting up to $L$ preceding query frames (the available history in a real time deployment setting), without exceeding the matrix boundary: $\mathbf{C}_{q}\in\mathbb{R}^{N\times R}$, where $N=\min(L,q+1)$.
We apply z-score normalisation to $\mathbf{C}$ across columns and then across rows, to reduce local bias in the similarity submatrix. To compute the sequence-matched score at each reference index $j$, we calculate the trace of the most recent $k\times k$ (where $k=\min(N,j+1)$) diagonal block:
$\tilde{\mathbf{S}}[q,j]=\operatorname{trace}\left(\mathbf{C}_{[N-k:N,\;j-k+1:j+1]}\right).$ (6)
@@ III-E Ensembling Strategies
Each $\tilde{\mathbf{S}}^{(m)}\in\mathbb{R}^{Q\times R}$ is the sequence-matched similarity matrix produced by the $m$-th method configuration. We treat these as individual ensemble members: $\mathcal{S}=\left\{\tilde{\mathbf{S}}^{(1)},\ \dots,\ \tilde{\mathbf{S}}^{(M)}\right\}$.
To aggregate predictions across configurations, we perform element-wise summation of all aligned similarity matrices, followed by maximum similarity selection as per the ensembling strategy from [11]:
@@ IV-A Datasets
In this paper, we evaluate our method on two outdoor datasets: BrisbaneEventVPR[11] and NSAVP[23], recorded during real-world vehicle traversals across day and night conditions.
BrisbaneEventVPR is a widely used benchmark dataset for event-based visual place recognition, introduced in 2020 and evaluated in several works [13, 10, 12, 8]. It consists of six traverses of an 8km route in Brisbane, Australia, recorded under various weather and illumination conditions, including daytime, night, sunrise, sunset, and morning on single and multilane roads. Data was captured using a $346\times 260$ resolution iniVation DAVIS 346 event camera mounted on the vehicle’s windshield, along with GPS and 1080p RGB video from a consumer-grade camera.
Novel Sensors for Autonomous Vehicle Perception (NSAVP) is a more recent dataset collected in Michigan, USA, with a focus on evaluating modern sensor technologies in diverse and realistic environments. It comprises two distinct routes (8.3 km and 8.6 km) traversed in both directions under varying lighting conditions. Data was collected using a $1280\times 720$ resolution Prophesee Gen4 HD event sensor and high-precision 200 Hz ground-truth from an Applanix POS-LV 420 system, yielding a higher-resolution, roof-mounted configuration from a different manufacturer than the DAVIS 346 used in BrisbaneEventVPR. Compared to BrisbaneEventVPR, NSAVP also introduces greater environmental diversity, including suburban neighbourhoods and dense urban scenes with multi-story buildings, and offers higher fidelity sensing for large-scale benchmarking.
To ensure fair comparison across illumination changes, we evaluate five day-to-day and five day-to-night pairs. On BrisbaneEventVPR [11], we use sunset1 as the reference against sunrise, morning, daytime, sunset2, and night. On NSAVP [23], we adopt the notation R0-[F/R]-[A/S/N] (Route 0, Forward/Reverse, Afternoon/Sunset/Night) to evaluate five pairs: FA0 vs. FS0, FA0 vs. FN0, FN0 vs. FS0, RA0 vs. RN0, and RS0 vs. RN0. Datasets such as DDD20 [40] and NYC Events [41] were excluded as they lack explicit repeat traversals.